You wish to test the following claim (HaHa) at a significance level of α=0.005α=0.005. For the context of this problem, one data set represents a pre-test and the other data set represents a post-test.     

Ho:μd=0Ho:μd=0
Ha:μd≠0Ha:μd≠0

You believe the population of difference scores is normally distributed, but you do not know the standard deviation. You obtain pre-test and post-test samples for n=329n=329 subjects. The average difference (post - pre) is ¯d=−1.8d¯=-1.8 with a standard deviation of the differences of sd=32.9sd=32.9.

  1. What is the test statistic for this sample?

    test statistic =   Round to 4 decimal places.
  2. What is the p-value for this sample? Round to 4 decimal places.

    p-value =  
  3. The p-value is...
    • less than (or equal to) αα
    • greater than αα

  4. This test statistic leads to a decision to...
    • reject the null
    • accept the null
    • fail to reject the null

  5. As such, the final conclusion is that...
    • There is sufficient evidence to warrant rejection of the claim that the mean difference of post-test from pre-test is not equal to 0.
    • There is not sufficient evidence to warrant rejection of the claim that the mean difference of post-test from pre-test is not equal to 0.
    • The sample data support the claim that the mean difference of post-test from pre-test is not equal to 0.
    • There is not sufficient sample evidence to support the claim that the mean difference of post-test from pre-test is not equal to 0.
Answer #1

given data are

sample n=329

average difference =-1.8

standard deviation of the difference sd=32.9

null hypothesis H0: - =0

alternative hypothesis Ha :-

= level of significance =0.005

test statistic :- t=

t= -1.8/(32.9/)

= -1.8 / 1.8138

=-0.9924

at df= n-1 =328 in 2 tail p value is

2× p (t= -0.9924)=2 ×0.16087

= 0.32174 = 0.3217

decision:- the p value is greater than so fail to reject the null hypothesis H0 .

conclusion:- There is not sufficient  sample evidence to support the claim that the mean difference of post- test from pre- test is not equal to 0